### v1

# Orbital Magnetization Reveals Multiband Topology

Abstract
We demonstrate that nontrivial multiband topological invariants of electronic wavefunctions can be revealed through diamagnetic orbital magnetization responses to external magnetic fields. We find that decomposing orbital magnetization into energetic and quantum-geometric contributions allows one to deduce nontrivial multiband topology, provided knowledge of the energy spectrum. We showcase our findings in general effective models with multiband Euler topology. We moreover identify such multiband topological invariants in effective models of strontium ruthenide ($\text{Sr}_{2}\text{Ru}\text{O}_{4}$), which may in principle be verified in the state-of-the-art doping-dependent magnetization measurements. Our reconstruction scheme for multiband invariants sheds a topological perspective on the multiorbital effects in materials realizing unconventional phenomenologies of orbital currents or multiband superconductivity.
Introduction.— Multistate effects play an important role in physics, from nonlinear optics Boyd (2019); Jankowski and Slager (2024); Avdoshkin et al. (2025); Jankowski et al. (2025a), to diverse correlated quantum phenomena such as superconductivity Peotta and Törmä (2015); Xie et al. (2020); Peri et al. (2021); Herzog-Arbeitman et al. (2022a); Gassner et al. (2024); Chau et al. (2025) and magnetism Kang et al. (2024); Oh et al. (2024); Hu et al. (2025); g. Oh et al. (2025). In condensed matter systems, multistate effects are fundamentally encoded in the combinations of matrix elements of quantum operators between multiple single-particle band eigenstates Ahn et al. (2020, 2021); Bouhon et al. (2023); Yu et al. (2024); Jankowski and Slager (2024); Jankowski et al. (2025b); Avdoshkin et al. (2025); Jain et al. (2025). The underlying multiband geometry Bouhon et al. (2023); Törmä (2023); Jain et al. (2025) arises from the fact that any quantum operator is a derivative in a conjugate quantum variable Ahn et al. (2019); Shinada and Nagaosa (2025), which translates geometric tensors and differential structures of bands into quantum responses manifested in numerous experiments Lai et al. (2021); Yue et al. (2022); Wang et al. (2023); Gao et al. (2023). Combinations of quantum states can realize not only nontrivial Hilbert-space geometry Ahn et al. (2021); Bouhon et al. (2023); Yu et al. (2024), but also nontrivial multiband topologies that are irreducible to sums of single-band invariants such as Chern numbers in anomalous quantum Hall phases Niu et al. (1985); Haldane (1988); Yu et al. (2010) or topological crystalline invariants Fu (2011); Slager et al. (2012); Shiozaki and Sato (2014); Kruthoff et al. (2017); Po et al. (2017); Bradlyn et al. (2017).
Topologies associated with finite multiband invariants and multigap conditions Zhao and Lu (2017); Wu et al. (2019); Ahn et al. (2019); Bouhon et al. (2020a, b); Ünal et al. (2020); Guo et al. (2021); Jiang et al. (2021); Slager et al. (2024); Wahl et al. (2025) are an increasingly studied topic of interest, yet experimental probes of these exotic topologies are rather scarce. The connection to multiband geometry is of key importance for potential multiband invariant reconstruction schemes. For instance, nonlinear optical responses involving multiband torsion tensors allow to extract unconventional topological invariants beyond the tenfold classification Kitaev (2009); Davoyan et al. (2024); Jankowski and Slager (2024); Jankowski et al. (2024, 2025c). A longstanding challenge includes finding electromagnetic probes for multiband topological Euler invariants Jankowski et al. (2025b); Jain et al. (2025); Chau et al. (2025) in electronic materials, which at the time of writing, lack a general measurement scheme independent of band energies. In this direction, a particularly underexplored and promising direction entails potential novel responses to magnetic fields beyond the recently studied geometric responses of Euler bands to electric fields Jankowski et al. (2025b); Jain et al. (2025). Magnetic field probes are of particularly high promise for revealing nontrivial multiband topologies, as these (i) reflect the exotic Landau levels arising from multiband topologies Bouhon et al. (2020a); Herzog-Arbeitman et al. (2022b); Guan et al. (2022), and (ii) couple to angular momentum operators respecting multistate resolutions by definition Xiao et al. (2010). As such, the multistate topologies should be in principle anomalously manifested, and hence observable, through certain magnetic responses.
Figure 1: Orbital magnetization $(\textit{m})$ from diamagnetic loop currents $(j_{xy})$, in response to an out-of-plane magnetic field $(\textit{B})$. The diamagnetic orbital response fingerprints the nontrivial Euler band topology.
In this work, we identify the role of multistate topological invariants in orbital magnetic susceptibilities and diamagnetic responses. We find a rich interplay. First, we show that the orbital magnetic susceptibility can be naturally decomposed in terms of energetic and quantum geometric contributions Gao et al. (2015); Ogata and Fukuyama (2015). The former can be estimated directly from the energy spectrum and fermionic occupation numbers determined by the quasiparticle energies and chemical potential. Thus, the energetic contribution can be deduced solely from the angle-resolved photoemission spectroscopy (ARPES) experiments. On the other hand, we find that the geometric term explicitly involves the quantum metric tensor and encodes the information about the topological invariants realized by electronic states in the crystalline material. Second, we leverage this interplay of energetic and geometric orbital magnetic susceptibility terms to reconstruct multiband topological invariants in continuum and lattice models with multiband Euler topology. Third, we showcase the correspondence of orbital magnetic susceptibility and multiband topological invariant in a material realization context of strontium ruthenide ($\text{Sr}_{2}\text{Ru}\text{O}_{4}$) models. Therefore, our findings not only demonstrate that quantum geometry plays an important role in orbital magnetic susceptibility, but also allows one to experimentally reveal the presence of exotic multiband topological invariants realized in quasiparticle bands. Hence, we establish the orbital magnetic susceptibility as an observable of interest for probing multiband topologies, analogously to the anomalous Hall conductance constituting the hallmark observable for single-band Chern invariants in topological matter. In the following, we detail how the orbital magnetic susceptibility arises as a probe of multiband topology, and captures the multiband quantum geometry even in effective models of real materials.
Results.— We now demonstrate how the multiband topology can be revealed via orbital magnetization measurements in diamagnetic responses of quasi-two-dimensional crystals. The orbital magnetization $m$ is given by an orbital susceptibility $\chi_{\text{O}}$, $m=\chi_{\text{O}}B$, upon applying an external out-of-plane magnetic field $B$ (see Fig. 1). In terms of current operators $j_{x},j_{y}$, the orbital magnetic susceptibility $\chi_{\text{O}}$ is given by Fukuyama (1971):
$\chi_{\text{O}}=\frac{e^{2}}{\hbar^{2}c^{2}}k_{B}T\sum_{n,{\bf k}}\mathrm{Tr}[Gj_{x}Gj_{y}Gj_{x}Gj_{y}]\equiv\frac{e^{2}}{\hbar^{2}c^{2}}k_{B}T\sum_{n,{\bf k}}\mathrm{Tr}[j_{xy}j_{xy}]\;,$ (1)
where we sum over Matsubara frequencies $n$ and single-particle momenta k. We introduce a shorthand notation $j_{xy}=Gj_{x}Gj_{y}$ in terms of the electronic Green’s function, propagator $G$. Centrally to this work, we recast the orbital magnetization in terms of gauge-independent energetic ($\chi_{\text{E}}$) and geometric contributions ($\chi_{\text{geo}}$), $\chi_{\text{O}}=\chi_{\text{E}}+\chi_{\text{geo}}$. The energetic orbital susceptibility term $\chi_{\text{E}}$ in the decomposition reads SI ,
$\begin{aligned} \chi_{\text{E}}&=\frac{e^{2}}{6\hbar^{2}c^{2}}\sum_{a}\sum_{\bf k}\left[2\left(\frac{\partial^{2}\epsilon_{a}}{\partial k_{x}\partial k_{y}}\right)^{2}+\frac{\partial^{2}\epsilon_{a}}{\partial k_{x}^{2}}\frac{\partial^{2}\epsilon_{a}}{\partial k_{y}^{2}}\right.\\ &\left.+\frac{3}{2}\left(\frac{\partial\epsilon_{a}}{\partial k_{x}}\frac{\partial^{3}\epsilon_{a}}{\partial k_{x}\partial k_{y}^{2}}+\frac{\partial\epsilon_{a}}{\partial k_{y}}\frac{\partial^{3}\epsilon_{a}}{\partial k_{x}^{2}\partial k_{y}}\right)\right]f^{\prime}_{a}\end{aligned}\;,$ (2)
which can be directly estimated from the energy spectrum ($\epsilon_{a}$) measurable in ARPES experiments. $a$ labels the quasiparticle bands, and $f_{a}(\mu)=\big[\text{exp}(\frac{\epsilon_{a}-\mu}{k_{B}T})+1\big]^{-1}$ is the Fermi-Dirac distribution function, which establishes equilibrium occupations under the known chemical potential $\mu$. The quantum geometric contribution to orbital susceptibility Piéchon et al. (2016), $\chi_{\mathrm{geo}}$, can be decomposed in terms of three geometric terms, $\chi_{\mathrm{geo}}=\chi_{xx}+\chi_{yy}+\chi_{xy}$}, which read
$\chi_{xx}$ $=\frac{e^{2}}{\hbar^{2}c^{2}}\sum_{a\neq b}\sum_{\bf k}\partial_{y}\epsilon_{a}\partial_{y}\epsilon_{b}g^{ab}_{xx}\left(f^{\prime}_{a}+f^{\prime}_{b}-2\frac{f_{ab}}{\epsilon_{ab}}\;\right),$ (3)
$\chi_{yy}$ $=\frac{e^{2}}{\hbar^{2}c^{2}}\sum_{a\neq b}\sum_{\bf k}\partial_{x}\epsilon_{a}\partial_{x}\epsilon_{b}g^{ab}_{yy}\left(f^{\prime}_{a}+f^{\prime}_{b}-2\frac{f_{ab}}{\epsilon_{ab}}\right)\;,$ (4)
$\chi_{xy}$ $=4\frac{e^{2}}{\hbar^{2}c^{2}}\sum_{a\neq b}\sum_{\bf k}\frac{1}{\epsilon_{ab}}\partial_{x}\epsilon_{b}\partial_{y}\epsilon_{b}g^{ab}_{xy}\left(f_{ab}-\epsilon_{ab}f_{a}^{\prime}+\frac{1}{2}\epsilon_{ab}^{2}f_{a}^{\prime\prime}\right)\;.$ (5)
[…]
Figure 2: Topological bands with Euler charge $e=1$ in Lieb lattice model. (a) Low-energy bands supporting the Euler invariant via a quadratic band touching. (b) The quantum geometric contribution to orbital magnetic susceptibility associated with the Euler charge (Geo) dominates the band-dispersive susceptibility contribution ($\text{Eng})$. The dashed line indicates the energy of the band touching with nontrivial Euler class.
We now showcase how the geometric orbital magnetic susceptibility term $\chi_{\text{geo}}$ allows to deduce the Euler class invariant $e\in\mathbb{Z}$ Bouhon et al. (2019); Ahn et al. (2019). The Euler invariant $e$ is defined using the multiband connection between bands $a$ and $b$ expressed in a real gauge, ${A}_{\mu}^{ab}=\langle u_{a{\bf k}}|\partial_{\mu}u_{b{\bf k}}\rangle$, which determines the Euler form $\mathrm{Eu}^{ab}=\partial_{x}{A}_{y}^{ab}-\partial_{y}{A}_{x}^{ab}$} Ahn et al. (2019); Bouhon et al. (2020a). The multiband Euler invariant over a two-dimensional Brillouin zone (BZ) patch $\mathcal{D}$ reads Ahn et al. (2019),
$e=\frac{1}{2\pi}\left[\int_{\mathcal{D}}\text{d}^{2}{\bf k}~\mathrm{Eu}^{ab}-\oint_{\partial\mathcal{D}}\text{d}{\bf k}\cdot{\textbf{A}}^{ab}\right]\;\in\mathbb{Z},$ (6)
and captures an obstruction to Stokes’s theorem applied to the multiband connection. The quantization of the invariant is protected by a spatiotemporal inversion ($\mathcal{PT}$) symmetry Bouhon et al. (2019); Ahn et al. (2019). The minimal integer Euler invariant, $|e|=1$, can be realized by a quadratic band touching within a general class of ${\bf k}\cdot\mathbf{p}$ models expressed in terms of band projector matrices $P_{1},P_{2}$, see also SM SI ,
$H_{\text{Eu}}$ $=\frac{k^{2}}{2m_{1}}P_{1}+\frac{k^{2}}{2m_{2}}P_{2}=\frac{1}{4m_{1}m_{2}}[(m_{1}+m_{2})k^{2}\sigma_{0}$
$+2(m_{1}-m_{2})k_{x}k_{y}\sigma_{1}+(m_{1}-m_{2})(k_{x}^{2}-k_{y}^{2})\sigma_{3}]\ .$ (7)
$m_{1}$ and $m_{2}$ are the effective masses of the touching bands, $k^{2}=k_{x}^{2}+k^{2}_{y}$, and $\sigma_{i}$ are the Pauli matrices. In Fig. 2(a), we present a quadratic band touching realizing an Euler class charge $|e|=1$ in bands with effective masses $m_{1},m_{2}<0$. In Fig. 2(b), we demonstrate the corresponding orbital magnetic susceptibility $\chi_{\text{O}}$, decomposed in terms of $\chi_{\text{E}}$ and $\chi_{\text{geo}}$ contributed by the multiband quantum metric $g_{\mu\nu}^{ab}=\text{Re}({A}_{\mu}^{ab}{A}_{\nu}^{ba})$. For full analytical derivation, see SM SI . We note that the energetic contribution $(\chi_{\text{E}})$ is order of magnitude smaller than the topologically-induced geometric contribution $(\chi_{\text{geo}})$ to the total orbital magnetic susceptibility $(\chi_{\text{O}})$.
Furthermore, beyond the general effective ${\bf k}\cdot\mathbf{p}$ expansions, we demonstrate an emergence of analogous multiband topological quantum geometric susceptibility contributions in models of strontium ruthenide ($\text{Sr}_{2}\text{Ru}\text{O}_{4}$Noce and Cuoco (1999). Correspondingly, we employ a tight-binding Hamiltonian Noce and Cuoco (1999) on a Lieb lattice (see SM SI ) to fit the theoretical band energies to the experimentally accessible quasiparticle energies measured in ARPES. In Fig. 3, we show the electronic bands within a seven-band model, where we retrieve the Euler invariants $|e|=1$ in the higher energy band subspace close to the Fermi level. Figure 3(b) demonstrates the geometric orbital susceptibility contributions contributed by the Euler charges in distinct band subspaces. Once again, we find that $\chi_{\text{O}}$ depends heavily on the topologically-induced singular contribution of quantum metric $g_{\mu\nu}^{ab}$, which is realized between distinct pairs of bands contributing to $\chi_{\text{geo}}$. The remaining enhanced contribution to orbital susceptibility in Fig. 3(b) arises from van Hove (vH) singularities, rather than being a multiband topological effect. This results in a combination of enhanced density of states and the corresponding occupation factors. However, the vH singularity contributions can be disregarded as nontopological purely from the inspection of ARPES spectra. Thus, upon disregarding the expected vH contributions, $\chi_{\text{O}}$ allows reconstructing the nontrivial multistate geometry and multiband topologies purely from $\chi_{\text{geo}}$ and its dependence on doping ($\mu$), as we address next.
Discussion and conclusions.— We further discuss how the multiband topological invariant can be deduced from the diamagnetic orbital magnetization response. Moreover, we provide a semiclassical interpretation of our findings, which we phrase in terms of the multiband wavepacket dynamics.
Figure 3: Orbital magnetic susceptibility fingerprints topological Euler invariants in Sr2RuO4 model. (a) Band structure obtained from a tight-binding model of Sr2RuO4. Bands contributed by the $(\mathrm{Ru}\;d_{xz},\;\mathrm{Ru}\;d_{yz},\;\mathrm{O}1\;p_{z},\;\mathrm{O}2\;p_{z})$ orbitals (solid lines) realize the nontrivial Euler class invariants, $|e|=1$ $(\text{Eu}_{1,2})$, and van Hove singularities (vH); dotted lines mark $(\mathrm{Ru}\;d_{xy},\;\mathrm{O}1\;p_{x},\;\mathrm{O}2\;p_{y})$ orbital bands. (b) Orbital magnetic susceptibility $\chi$ decomposed into energetic (Eng) and geometric (Geo) contributions as a function of chemical potential $\mu$, with $\chi_{0}=t_{0}(ea^{2}/4\pi^{2}\hbar c)^{2}$ where $a$ is the lattice constant and $t_{0}=1$ eV. Around $\mu=-0.8$ and $\mu=0.2~\text{eV}$, the total orbital magnetic susceptibility $\chi_{\text{O}}$ reflects the presence of topologically-induced concentrated quantum geometries due to the Euler charges.
To reconstruct the multiband invariant, one may first identify possible quantum geometry hotspots, such as band touchings observable in ARPES. In the case of trivial multiband invariants, only the energetic orbital susceptibility contribution $\chi_{\text{E}}$ arises. $\chi_{\text{E}}$ can be directly estimated based on the ARPES band structure data, from Eq. (2). We propose that the geometric contribution $\chi_{\text{geo}}$ can be thus extracted from combining ARPES and orbital magnetic susceptibility $\chi_{\text{O}}$ data with the $\chi_{\text{E}}$ estimation. Importantly, the known spin magnetic susceptibility contributions must be further accounted for, as the total magnetic susceptibility is measured g. Oh et al. (2025). The presence of a multiband topological invariant can be validated upon inserting an analytical ansatz for $g^{ab}_{\mu\nu}~\propto|e|^{2}/|k|^{2}$ in Eqs. (3-5), which follows from Eq. (7), in the momentum space neighborhood of the band touchings. The contribution arising from this ansatz can be compared against the experimental data.
It should be noted that although a resolution of the quantum metric tensor in momenta has been experimentally achieved in solid state materials very recently Kang et al. (2025); Kim et al. (2025), these findings are limited by a two-band approximation. In this work, we go beyond this paradigm. Namely, we demonstrate that the multiband metric and topological invariants can be inferred even in a seven-band model of $\text{Sr}_{2}\text{Ru}\text{O}_{4}$, assuming the full knowledge of the ARPES spectrum, as in the mentioned experiments.
Finally, we provide a semiclassical interpretation of our findings. The orbital susceptibility contributed by quantum geometry reflects diamagnetic currents, which semiclassically arise from a current loop model. Increasing magnetic field induces a Lorentz force on the wavepacket, which responds diamagnetically. The response arises from a field-induced positional shift correction in Berry connection, $A^{\prime}_{i}(B)=F_{ij}B_{j}$, with $F_{ij}$, the magnetic polarizability tensor introduced by Gao and Niu Gao et al. (2014, 2015). The $F_{ij}$ tensor involves the multiband connections, which connect the response directly to the Euler curvature form, and can support orbital magnetization via field-induced Berry curvature, $\Omega^{\prime}_{xy}(B)=\partial_{y}A^{\prime}_{x}(B)-\partial_{x}A^{\prime}_{y}(B)$. The associated positional wavepacket shift results in diamagnetic currents, as follows from a direct inspection. The further interplay of magnetic field-induced positional shifts and multiband geometries will be studied in a future work.
To sum up, we demonstrate that the presence of multiband Euler topology in a bulk material can be inferred from its diamagnetic orbital magnetization responses. Our findings establish orbital magnetization as an experimentally measurable physical property of choice for identifying the quantum geometric fingerprint of multiband topological Euler invariant. We thus propose an electromagnetic smoking-gun probe for pairs of topological bands supporting vorticity in a multiband Berry connection. Our work establishes a physical go-to response for probing Euler bands, analogously to the transverse Hall conductance being the hallmark response of Chern bands realizing intraband Berry curvature in quantum Hall phases.
Acknowledgments.— The authors thank Gaurav Chaudhary for discussions and comments on the manuscript. C.W.C. acknowledges funding from the Croucher Cambridge International Scholarship by the Croucher Foundation and the Cambridge Trust. R.-J.S. acknowledges funding from an EPSRC ERC underwrite Grant No. EP/X025829/1, and a Royal Society exchange Grant No. IES/R1/221060, as well as Trinity College, Cambridge. W.J.J. acknowledges funding from the Rod Smallwood Studentship at Trinity College, Cambridge.